OpenStudy (anonymous):

Pak Su rears some goats and chickens on his farm.If there are 6 eyes and 102 legs of animals on his farm,how many goats and chickens are there?

OpenStudy (anonymous):

6 eyes

OpenStudy (anonymous):

only 6 eyes? must be a bunch of blind animals on that farm >.< or maybe there are caterpillars and he didnt say anything about them o.O

OpenStudy (anonymous):

check if it is 6 eyes or more

OpenStudy (anonymous):

Assuming it's 60 eyes: First we setup the eyes: \[2x+2y=60\] Then we setup the legs: \[2x+4y=102\] Subtract the first problem from the second and get: \[2y=42\] \[y=21\] Now we substitute this back into the first problem: \[2x+2(21)=60\] \[2x=60-42\] \[2x=18\] \[x=9\] So the answer is 21 goats and 9 chickens, again assuming 60 eyes.